TRACE ATMOSPHERIC DETECTION OF HCHO VIA FIBER LASER-INDUCED FLUORESCENCE DUR-ING BEARPEX 2009

\underline{J. DIGANGI}, S. HENRY, A. KAMMRATH, E. BOYLE, and F. KEUTSCH, Department of Chemistry, University of Wisconsin-Madison, Madison, WI 53706.; J. PAUL, NovaWave Technologies, Redwood City, CA 94065.

Formaldehyde (HCHO) plays an important role in atmospheric chemistry as it is an oxidation product of volatile organic compounds (VOCs), a major source for HO_{x} radicals. We will present results from field measurements taken during Biosphere Effects on Aerosols and Photochemistry Experiment (BEARPEX) 2009 using FIber Laser-Induced Fluorescence (FILIF) detection of HCHO. Central to the FILIF technique is the first of a new class of UV fiber lasers from NovaWave Technologies. By using this narrow bandwidth (< 300 MHz) UV laser to excite a rotational feature in the 4_{0}^{1} A^{1}A_{2}←X^{1}A_{1} band at 353 nm, this technique is capable of very high HCHO sensitivity and selectivity . The combination of low power comsumption and weight, compact size, and ruggedness, provides an instrument ideally suited to field measurements.